OpenStudy (anonymous):

how to mx+b= y with two points (-2,0) (-2,-2)

OpenStudy (aroub):

first find the slope

OpenStudy (aroub):

and then pick any of the points u have!

OpenStudy (aroub):

like this : slope = -2-0/ -2+2 = -2/0 = undefined / u have an undefined slope!

OpenStudy (aroub):

so the answer is y=-2

OpenStudy (anonymous):

To define a line you need 2 points or a direction. To find the vector direction you just need to do for example (-2,0) - (-2,-2) = (0,2) This is the line equation formula: (x,y) = (-2,0) + k(0,2) if you solve for x you'll get: x = -2 + 0*k y = 0 +2*k (=) k = y/2 substituting on the "x" equation: x = -2 + 0 * (y/2) x = -2 this is a line parallel to y-axis... in y=mx + b format it will be: 0=1 * x +2

OpenStudy (aroub):

oh my answer is incorrect??

OpenStudy (anonymous):

You prolly made a mistake somewhere. Don't worry, it happens even to the best of the best :)

OpenStudy (aroub):

lol! oh yeaaa!! thank youu!! :)